William, Joseph, Margaret, Frank, and Jeffrey all work for the same company. Frank has been there the longest, eleven years and makes the most money, three times as much as Jeffrey. Margaret has been there the second longest and makes $50,000 a year. Joseph has worked there four years less than Frank and one year more than William. He makes $60,000 a year. William’s pay is twice as much as Margaret’s. Jeffrey has worked there three years less than William and his income is three-quarters of Margaret’s. How much is Frank's annual salary?
Answer #1
Just write down the given data

Margaret = 50000
Joseph = 60000
Jeffrey = (3/4)(50000) = 37500
Frank = 3(37500) = 112500